the weight of the object measured in the mars in 10 n where the observed gravitational acceleration is 3.40 m/ s^2 . determine the weight of the object in earth where the gravitational acceleration is 9. 81 m/s^2? solution


1
Expert's answer
2021-03-08T08:26:54-0500

"P=mg,\\P\sim g,"

"P_2=P_1\cdot \frac{g_2}{g_1},\\
P_2=10\cdot\frac{9.81}{3.40}=28.9~\text{N}."